OpenStudy (anonymous):

The length of a rectangle is 6 yards longer than it's width. If the perimeter of the rectangle is 32 yards , find it's area in square yards.

OpenStudy (anonymous):

We have to find the length first, so we set up an equation. if all the sides added together equal 32, and w+6=L, then our equation will be 4w+12=32. We subtract 12 from both sides to get 4w=20, and we divide both sides by 4 to get w=5. Now we know that the width is 5. To find the length we add 6 because the length is 6 yards longer then the width. L=6+5 L=11 To find the area in square yards, we multiply the length by the width 11*5=55 yards^2